# ImF: Implicit Fingerprint for Large Language Models

### v2

@@ Abstract.
journalyear: 2025copyright: acmlicensedconference: ; booktitle: ;ccs: Security and privacy Security servicesccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
Consequently, researchers have drawn considerable attention to the injected fingerprint, which can be verified without accessing the internal details of LLMs. The injected fingerprint generally embeds a unique Secretly pick$(x,y)$ within the model based on instruction poisoning and backdoor attacks, where $x,y$ represent the input and output of LLMs, respectively. Xu et al. (Xu et al., 2024) proposed the Instructional Fingerprinting (IF) method, which constructed the Secretly pick $(x,y)$ by adding unique markers such as scrambled text combining multiple languages and symbols, thereby preventing the fingerprint from being overwritten during downstream fine-tuning with similar training data. In contrast, Russinovich et al. (Russinovich and Salem, 2024) proposed the Chain &\&& Hash method, which aimed to enhance the invisibility of the fingerprint by using cryptographic techniques to construct the Secretly pick $(x,y)$ with a normal question and a specific word or a short phrase.
@@ 3. Erase the fingerprint by GRI attack
Due to the fingerprint output $y$ being designed based on special rules rather than the natural context of the input question $x$, explicitly Secretly pick $x,y$ has a weak semantic correlation. The Instructional Fingerprinting (IF) method (Xu et al., 2024) embeds explicit markers in the input $x$ (e.g., Garbled text that includes multiple languages and special characters) and maps to the output $y$ (e.g., the phrase ”This is Meta’s model”). Similarly, Chain &\&& Hash (C&H) (Russinovich and Salem, 2024) employs cryptographic techniques to select input-output pairs for fingerprint embedding. Specifically, C&H selects a set of normal questions as inputs $x$ and uses a hash function to determine corresponding outputs $y$ from a predefined answer bank containing simple phrases like ”One,” ”OK,” or similar short responses. However, this design results in $x$ and $y$ still maintaining weak semantics.
@@
9:        $bit$leftarrow determine embedding capacity based on $prob$ distribution
@@
11:        $int_{embed}\leftarrow$ read next $bit$ from $B$ starting at $bit_{index}$
12:        $prob$leftarrow select probability group corresponding to $int_{embed}$
13:        $bit_{index}$leftarrow $bit_{index}$ + $bit$
@@
15:     $w\leftarrow$ sample from $prob$
@@
24:  while $iteration\_count<T$ do
@@
[…]
27:     if similarity $>$ $\delta$ then
@@ 4.2. Stego Secretly pick $x$ Generation
$\text{IF}_{\text{Lora}}$ 30% 60% 70% 100% 70% 100% 100% 40%
@@
[…]
$\text{ImF}_{\text{Lora}}$ 100% 90% 100% 90% 100% 90% 70% 60%
@@
$\text{IF}_{\text{Lora}}$ 100% 100% 100% 100% 100% 100% 60%
@@
[…]
$\text{ImF}_{\text{Lora}}$ 100% 100% 100% 90% 100% 80% 100%
@@ 5.2.2. Robustness Analysis of Embedding Fingerprints for different scaling of LLMs
[…]
$\text{ImF}_{\text{SFT}}$ 100% 100% 100% 100% 100% 100% 100% 100%
@@
[…]
$\text{ImF}_{\text{SFT}}$ 100% 100% 100% 100% 100% 100% 100%
@@ 5.4. Harmlessness Analysis of Fingerprints
[…]
$\text{ImF}_{\text{Lora}}$ 100% 90% 100% 90% 100% 90% 70% 60%
@@
[…]
$\text{ImF}_{\text{Lora}}$ 100% 100% 100% 90% 100% 80% 100%

### v3

@@ Abstract.
journalyear: 2025copyright: acmlicensedconference: ; booktitle: ;$c$s: Security and privacy Security servicesccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
Consequently, researchers have drawn considerable attention to the injected fingerprint, which can be verified without accessing the internal details of LLMs. The injected fingerprint generally embeds a unique Secretly pick(x,y)(x,y) within the model based on instruction poisoning and backdoor attacks, where $x,y$ represent the input and output of LLMs, respectively. Xu et al. (Xu et al., 2024) proposed the Instructional Fingerprinting (IF) method, which constructed the Secretly pick $(x,y)$ by adding unique markers such as scrambled text combining multiple languages and symbols, thereby preventing the fingerprint from being overwritten during downstream fine-tuning with similar training data. In contrast, Russinovich et al. (Russinovich and Salem, 2024) proposed the Chain &\& Hash method, which aimed to enhance the invisibility of the fingerprint by using cryptographic techniques to construct the Secretly pick $(x,y)$ with a normal question and a specific word or a short phrase.
@@ 3. Erase the fingerprint by GRI attack
Due to the fingerprint output $y$ being designed based on special rules rather than the natural context of the input question $x$, explicitly Secretly pick $x,y$ has a weak semantic correlation. The Instructional Fingerprinting (IF) method (Xu et al., 2024) embeds explicit markers in the input $x$ (e.g., Garbled text that includes multiple languages and special characters) and maps to the output $y$ (e.g., the phrase ”This is Meta’s model”). Similarly, Chain &\& Hash (C&H) (Russinovich and Salem, 2024) employs cryptographic techniques to select input-output pairs for fingerprint embedding. Specifically, C&H selects a set of normal questions as inputs $x$ and uses a hash function to determine corresponding outputs $y$ from a predefined answer bank containing simple phrases like ”One,” ”OK,” or similar short responses. However, this design results in $x$ and $y$ still maintaining weak semantics.
@@
9:   bitbit $\leftarrow$ determine embedding capacity based on probprob distribution
@@
11:   $int_{embed}\leftarrow$ read next bitbit from $B$ starting at $bit_{index}$
12:   probprob $\leftarrow$ select probability group corresponding to $int_{embed}$
13:   $bit_{index}$ $\leftarrow$ $_{index} + bitbit$
@@
15:  $w\leftarrow$ sample from probprob
@@
24:while iteration_count<Titeration\_count<T do
@@
26:  similarity $\leftarrow\text{ComputeSimilarity}(y,y_{1})$
27:  if similarity >> $\delta$ then
@@ 4.2. Stego Secretly pick $x$ Generation
$\text{IF}_{\text{Lora}}$ - 30% 60% 70% 100% 70% 100% 100% 40%
@@
$\text{C\&H}_{\text{Lora}}$ - 80% 10% 100% 70% 100% 70% 20% 10%
$\text{C\&H}_{\text{Lora}}$ GRI 40% 0% 40% 10% 20% 10% 0% 0%
$\text{ImF}_{\text{Lora}}$ - 100% 90% 100% 90% 100% 90% 70% 60%
@@
$\text{IF}_{\text{Lora}}$ - 100% 100% 100% 100% 100% 100% 60%
@@
$\text{C\&H}_{\text{Lora}}$ - 50% 100% 30% 80% 100% 0% 40%
$\text{C\&H}_{\text{Lora}}$ GRI 30% 50% 20% 50% 100% 0% 0%
$\text{ImF}_{\text{Lora}}$ - 100% 100% 100% 90% 100% 80% 100%
@@ 5.2.2. Robustness Analysis of Embedding Fingerprints for different scaling of LLMs
$\text{C\&H}_{\text{SFT}}$ - 100% 100% 100% 100% 100% 100% 100% 90%
$\text{C\&H}_{\text{SFT}}$ ft 100% 100% 100% 100% 100% 90% 20% 10%
$\text{C\&H}_{\text{SFT}}$ GRI 100% 100% 100% 100% 100% 100% 100% 80%
$\text{C\&H}_{\text{SFT}}$ ft+GRI 20% 90% 90% 70% 100% 70% 20% 0%
$\text{C\&H}_{\text{SFT}}$ merge 0% 100% 0% 100% 100% 100% 100% 0%
$\text{ImF}_{\text{SFT}}$ - 100% 100% 100% 100% 100% 100% 100% 100%
@@
$\text{C\&H}_{\text{SFT}}$ - 100% 100% 100% 100% 100% 100% 100%
$\text{C\&H}_{\text{SFT}}$ ft 100% 50% 90% 70% 100% 80% 10%
$\text{C\&H}_{\text{SFT}}$ GRI 100% 100% 100% 100% 100% 100% 100%
$\text{C\&H}_{\text{SFT}}$ ft+GRI 60% 20% 90% 50% 90% 40% 10%
$\text{C\&H}_{\text{SFT}}$ merge 100% 0% 100% 100% 100% 100% 100%
$\text{ImF}_{\text{SFT}}$ - 100% 100% 100% 100% 100% 100% 100%
@@ 5.4. Harmlessness Analysis of Fingerprints
$\text{ImF-w'CoT'}_{\text{Lora}}$ - 30% 90% 40% 100% 90% 90% 0% 0%
$\text{ImF-w'CoT'}_{\text{Lora}}$ GRI 10% 70% 0% 30% 60% 30% 0% 0%
$\text{ImF-w'CoT'}_{\text{SFT}}$ - 70% 70% 70% 90% 100% 90% 70% 70%
$\text{ImF-w'CoT'}_{\text{SFT}}$ GRI 70% 70% 70% 90% 100% 100% 70% 60%
$\text{ImF}_{\text{Lora}}$ - 100% 90% 100% 90% 100% 90% 70% 60%
@@
$\text{ImF-w'CoT'}_{\text{Lora}}$ - 100% 90% 90% 60% 100% 30% 10%
$\text{ImF-w'CoT'}_{\text{Lora}}$ GRI 80% 20% 70% 20% 90% 0% 0%
$\text{ImF-w'CoT'}_{\text{SFT}}$ - 100% 100% 80% 80% 100% 70% 70%
$\text{ImF-w'CoT'}_{\text{SFT}}$ GRI 100% 100% 80% 80% 100% 70% 70%
$\text{ImF}_{\text{Lora}}$ - 100% 100% 100% 90% 100% 80% 100%